Let me start by saying I am 37 and have taken on homeschooling for my 7th and 9th grader. My dilemma is I have searched everywhere for the answer to this problem. First of all I assume it is suppose to be searching for square feet, but it seems to be miss worded. So I'm confused. So here is the word problem verbatim... GEOMETRY To find the area of a floor, you can multiply its length times its width. The measure of the area of a floor is 49 units3. Find the most likely length and with of the room... So could you please help me understand it? My daughter worked it as 7x7=49. Would that be correct? It seemed to simple to be correct considering its geometry. Thank you so much, Tamela
Answer by fcabanski(1391)

Area is in square feet. The formula is length * width.


Your daughter's answer, a square room with sides = 7 (7 * 7 = 49) is correct.


The room could be 5 by 9.8. 5*9.8 = 49. But it's odd for rooms to have such measurements.


Most geometry is simple when you know the formulas.